# Training PINNs with Hard Constraints and Adaptive Weights: An Ablation Study

### v2

@@ 3.1 Allen-Cahn Equation $I$
The Allen-Cahn type PDE is a classical phase-field model that has been widely used to study phase separation phenomena bazant2017thermodynamic . It has numerous practical applications across a range of fields, including materials science allen1979microscopic ; shen2010numerical , biological systems hyman2014liquid ; takatori2015towards , and electrochemical systems horstmann2013precipitation ; tian2015electrochemically . We begin with the one-dimensional Allen-Cahn equation with periodic boundary conditions, a representative example that presents challenges for traditional PINN models, formulated as follows:
$u_{t}-\gamma_{1}u_{xx}+\gamma_{2}u^{3}-\gamma_{2}u=0,\quad(t,x)\in(0,T)\times(a,b),$ (3)
$u(0,x)=u_{0}(x),\quad x\in[a,b],$
$u(t,a)=u(t,b),\quad t\in[0,T],$
$u_{x}(t,a)=u_{x}(t,b),\quad t\in[0,T],$
@@
where $\gamma_{1},\gamma_{2}>0,T>0,a<b$ are prescribed constants. As $\gamma_{2}$ increases, the transition interface of the solutions is sharper, which makes it harder to solve the AC equation numerically. We first demonstrate the effectiveness of our scheme by testing on Allen-Cahn equation with the following initial condition and parameters, $u_{0}(x)=x^{2}\cos(\pi x)$, $T=1$, $a=-1$, $b=1$, $\gamma_{1}=0.001$, and $\gamma_{2}=5$.
Figure 1: Exact solution of the 1D Allen-Cahn Equation (Case I) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $1.18\times 10^{-3}$ and the relative $L^{1}$ error is $5.34\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 2: Best solutions of the Allen-Cahn Equation (Case I) $\gamma_{2}=5$. The resulting relative $L^{2}$ error is $1.18\times 10^{-3}$ and the relative $L^{1}$ error is $5.34\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
@@
For this example, we first test the baseline PINN approach raissi2019physics . As shown in Table 1, the baseline PINN alone fails to accurately solve the Allen-Cahn equation, motivating the development of a more effective training pipeline. To address this, we employ all proposed techniques (except mini-batching) and conduct an ablation study by disabling individual components one at a time. Additionally, we introduce two different decay functions for the initial condition, $e^{-1t}$ and $e^{-0.1t}$, to investigate their influence on the solution. The results are summarized in Table 1. The full scheme, including self-adaptive loss balancing, achieves the best performance, reaching a relative $L^{2}$ error of $1.18\times 10^{-3}$ and relative $L^{1}$ error of $5.34\times 10^{-4}$. In particular, using the faster-decaying initial condition ($IC1$) leads to better results than the slower-decaying case ($IC2$), suggesting that the initial condition plays a more critical role during the early stages of training but becomes less dominant over time. The ablation results further highlight the importance of each methodological component. Removing any single part generally leads to performance degradation, with the most severe impact observed when the initial condition is omitted, resulting in a relative error close to $0.99$ in both $L^{1}$ and $L^{2}$ norms. Figure 1 illustrates the exact solution, the network prediction, and the absolute error distribution. As seen, the predicted solution matches the ground truth remarkably well, with errors concentrated at negligible levels. Figure 2 presents cross-sectional comparisons at different times, further confirming the excellent agreement between the predicted and true solutions across the entire time horizon. These results provide strong evidence for the effectiveness of the proposed training strategy and underscore the critical role of each pipeline component.
Table 1: Allen-Cahn Equation (Casse $I$ and Case $II$): Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance - Case $I$ Performance - Case $II$
@@
$1.18\times 10^{-3}$ $5.34\times 10^{-4}$ $3.53\times 10^{-4}$ $2.14\times 10^{-4}$
@@
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
@@
$9.99\times 10^{-1}$ $9.98\times 10^{-1}$ $5.11\times 10^{-1}$ $3.28\times 10^{-1}$
3.2 Allen-Chan, Case $II$
For our second example, we investigate the performance of PINNs in scenarios involving sharper moving interfaces. Specifically, we consider a modified Allen–Cahn equation with more oscillatory initial data:
$u_{t}-\gamma_{1}u_{xx}+\gamma_{2}u^{3}-\gamma_{2}u=0,\quad(t,x)\in(0,1)\times(-1,1),$ (4)
$u(0,x)=x^{2}\sin(2\pi x),\quad x\in[-1,1],$
$u(t,-1)=u(t,1),\quad t\in[0,1],$
$u_{x}(t,-1)=u_{x}(t,1),\quad t\in[0,1],$
where $\gamma_{1}=0.001,\gamma_{2}=4$.
@@ 3.2 Allen-Chan, Case $II$
Figure 3: Exact solution of the Allen-Cahn Equation (Case II) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 4: Best solutions of the Allen-Cahn Equation (Case II) $\gamma_{2}=4$. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
3.3 Cahn-Hilliard Equation
Figure 5: Exact solution of the 1D Cahn-Hilliard Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $9.75\times 10^{-4}$ and the relative $L^{1}$ error is $5.65\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 6: Best solutions of the 1D Cahn-Hilliard Equation. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$ and the relative $L^{1}$ error is $2.14\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
The Cahn-Hilliard equation, a variant of the Allen-Cahn equation, also models phase separation. This process describes how the components of a binary fluid spontaneously separate into distinct domains, each composed of a single component cahn1958free ; kim2016basic ; lee2014physical ; miranville2017cahn . The equation is expressed as:
$u(0,x)=u_{0}(x),\quad x\in[-L,L],$
$u(t,-L)=u(t,L),\quad t\in[0,T],$
For the ablation study in Section 2, the neural network is trained using the same hyperparameters outlined in Section 3. In this experiment, the full scheme yields the best results. Additionally, methods using initial condition IC2 outperform those with IC1. Except for the full schemes and the experiment with periodic neural networks for boundary conditions and initial condition IC2, none of the other approaches produce accurate solutions. The errors for all experiments are summarized in Table 2. The best results are illustrated in Figures 5 and 6.
@@ 3.3 Cahn-Hilliard Equation
PBC IC1 (C=1) IC2 (C=0.1) SA Rel. $L^{2}$ error Rel. $L^{1}$ error
$1.06\times 10^{-3}$ $5.61\times 10^{-4}$
$9.75\times 10^{-4}$ $5.65\times 10^{-4}$
$2.50\times 10^{-1}$ $5.87\times 10^{-2}$
$3.87\times 10^{-3}$ $1.91\times 10^{-3}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$3.48\times 10^{-1}$ $2.94\times 10^{-1}$
$2.37\times 10^{-1}$ $1.09\times 10^{-1}$
$4.07\times 10^{-1}$ $3.45\times 10^{-1}$
$1.55\times 10^{-2}$ $8.20\times 10^{-3}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$9.99\times 10^{-1}$ $9.98\times 10^{-1}$
3.4 Kuramoto-Sivashinsky Equation
Kuramoto–Sivashinsky equation is a fourth-order nonlinear partial differential equation, known for its chaotic behavior kudryashov1990exact ; michelson1986steady . It has served as a pivotal model in the study of complex dynamical phenomena observed in various physical systems. Consequently, the Kuramoto–Sivashinsky equation provides an excellent case study to demonstrate how PINNs can effectively simulate chaotic dynamics. The equation is given by:
$u_{t}=-u_{xx}-u_{xxxx}-uu_{x},\quad(t,x)\in(0,T]\times(a,b),$ (6)
$u(0,x)=u_{0}(x),\quad x\in[a,b],$
$u(t,a)=u(t,b),\quad t\in[0,T],$
where $T=20,a=0,b=32\pi$, and $u_{0}(x)=\cos\frac{x}{16}(1+\sin\frac{x-1}{16})$. The parameters are the same as those mentioned in Section 3. The best results are shown in Figure 7 and 8:
Figure 7: Exact solution of the 1D Kuramoto-Sivashinsky Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $1.46\times 10^{-3}$ and the relative $L^{1}$ error is $1.02\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 8: Best solutions of the 1D Kuramoto-Sivashinsky Equation. The resulting relative $L^{2}$ error is $1.46\times 10^{-3}$ and the relative $L^{1}$ error is $1.02\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Table 3: Kuramoto-Sivashinsky Equation: Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance
PBC IC1 (C=1) IC2 (C=0.1) SA Rel. $L^{2}$ error Rel. $L^{1}$ error
@@ 3.4 Kuramoto-Sivashinsky Equation
As in Table 3, solving this PDE system using a baseline PINN alone proves to be exceptionally challenging. The incorporation of periodic neural networks for boundary conditions, paired with initial condition IC2, consistently yields superior results compared to IC1. This suggests that, in contrast to the Allen-Cahn equation discussed in Section 3.1, the choice of initial condition plays a more significant role in the training process for this more complex system. This observation underscores the importance of carefully selecting both the boundary conditions and initial conditions when training neural networks for chaotic or higher-order PDE systems, where these factors contribute notably to the accuracy and stability of the solution.
3.5 Gray-Scott Equation
Reaction and diffusion of chemical species can produce a variety of patterns de1991turing ; nishiura1999skeleton , reminiscent of those often seen in nature. The Gray-Scott type system is one of classical mathematical models for chemical reactions gray1983autocatalytic ; gray1984autocatalytic ; liang2022reversible . The general irreversible Gray-Scott equations describe such reactions:
@@ 3.5 Gray-Scott Equation
$u_{t}=\epsilon_{1}u_{xx}+b(1-u)-uv^{2},\quad(t,x)\in(0,T]\times(-L,L),$ (8)
$v_{t}=\epsilon_{2}v_{xx}-(b+k)v+uv^{2},$
$u(0,x)=u_{0}(x),\quad v(0,x)=v_{0}(x),\quad\forall x\in[-L,L],$
$u(t,-L)=u(t,L),\quad v(t,-L)=v(t,L),\quad\forall t\in[0,T],$
where $T=20,L=50$, $\epsilon_{1}=1,\epsilon_{2}=0.01$ are diffusion rates, $b=0.02$ is the "feeding rate" that adds $U$, $k=0.0562$ is the "killing rate" that removes $V$. We set our initial conditions as:
$u_{0}(x)=1-\frac{\sin(\pi(x-50)/100)^{4}}{2},\quad v_{0}(x)=\frac{\sin(\pi(x-50)/100)^{4}}{4}.$
In this ablation study, we not only evaluate the performance of self-adaptive neural networks but also test our scheme with mini-batching on the Gray-Scott equations, which model the densities of two interacting species. For mini-batching, we use 5 batches, each containing 3,280 samples. As shown in Table 4, the results confirm that each proposed component improves the overall model performance. Omitting any one of these components leads to higher error rates. Notably, the errors for the $u-$equation are typically smaller than those for the $v-$equation, which aligns with our expectations. This is because the $v-$equation is sharper and more challenging to train compared to the $u-$equation. Additionally, experiments using initial condition IC2 consistently outperform those using IC1. Specifically, the full scheme with IC2 achieves the best results, with a relative $L^{2}$ error of $2.05\times 10^{-4}$ for $u-$equation and $8.46\times 10^{-4}$ for $v-$equation. These results suggest that IC2 is more effective for larger time-scale PDEs, as IC1 causes the initial condition to decay too quickly, reducing its influence during training. The predicted solutions from our best model are visualized in Figures 9 and 10, respectively.
Figure 9: Best solutions of the 1D Gray-Scott Equation (u species). The resulting relative $L^{2}$ error is $2.05\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 10: Best solutions of the 1D Gray-Scott Equation (v species). The resulting relative $L^{2}$ error is $2.05\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 11 and Figure 12 show the comparison of the best prediction against the reference solution. Note that the majority of this error is attributed to the sharpness of the PDEs.
Figure 11: Exact solution of the 1D Gray-Scott Eqaution (v specie) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $8.46\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
Figure 12: Exact solution of the 1D Gray-Scott Eqaution (v specie) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $8.46\times 10^{-4}$. The hyper-parameter configuration is specified in Section 3.
@@
PBC IC1 IC2 SA mini-batch Rel. $L^{2}$ error (u - eq.) Rel. $L^{2}$ error (v - eq.)
@@
3.6 Belousov-Zhabotinsky Equation
A Belousov–Zhabotinsky reaction is one of a class of reactions that serve as a classical example of non-equilibrium thermodynamics. It typically involves the oxidation of organic compounds by bromine in an acidic medium zhabotinsky2007belousov ; cassani2021belousov . The system of PDEs takes the following form:
$u_{t}=\epsilon_{1}u_{xx}+u+v-uv-u^{2},$ (9)
$v_{t}=\epsilon_{2}v_{xx}+w-v-uv,$
$w_{t}=\epsilon_{1}w_{xx}+u-w,\quad(t,x)\in(0,T]\times[-L,L]$
$u(0,x)=u_{0}(x),\quad v(0,x)=v_{0}(x),\quad w(0,x)=w_{0}(x),\quad\forall x\in[-L,L]$
$u(t,-L)=u(t,L),\quad v(t,-L)=v(t,L),\quad w(t,-L)=w(t,L),\quad\forall t\in[0,T],$
where $T=2,L=1$, $\epsilon_{1}=10^{-5},\epsilon_{2}=2\times 10^{-5}$ are diffusion rates. We let the initial conditions be:
$u_{0}(x)=\exp(-100(x+0.5)^{2}),\quad v_{0}(x)=\exp(-100x^{2}),\quad w_{0}(x)=\exp(-100(x-0.5)^{2}).$
The system of PDEs involving three species provides a valuable opportunity to evaluate our proposed schemes using the techniques described in Section 2 for simulating multispecies dynamics. The results, summarized in Table 5, show that the full scheme consistently achieves the best overall performance, in line with previous experiments. Interestingly, the errors observed for methods using IC1 and those using IC2 are not significantly different. Both IC1 and IC2, when used independently, are capable of producing satisfactory results. However, when periodic neural networks are used to enforce boundary conditions, the performance declines and becomes less stable, even though self-adaptive training and mini-batching provide some improvements. This observation suggests that incorporating boundary conditions introduces additional complexity into the training process. Further investigation into their effects would be a promising direction for future research.
Figure 13: Best solutions of the 1D Belousov-Zhabotinsky equation (u specie). The resulting relative $L^{2}$ error is $3.49\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 14: Best solutions of the 1D Belousov-Zhabotinsky equation (v specie). The resulting relative $L^{2}$ error is $2.45\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 15: Best solutions of the 1D Belousov-Zhabotinsky equation (w specie). The resulting relative $L^{2}$ error is $2.50\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
The comparisons between the exact solutions of $(u,v,w)$ equations with corresponding best neural network predictions on the whole $(x,t)$ domain are shown in Figures 16, 17 and 18. Although there is a slight misalignment between each two of them, the model predictions closely match the corresponding numerical estimations.
Figure 16: Exact solution of the 1D Belousov-Zhabotinsky equation (u specie) with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $3.49\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 17: Exact solution of the 1D Belousov-Zhabotinsky equation (v specie) Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $2.45\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Figure 18: Exact solution of the 1D Belousov-Zhabotinsky equation (w specie) Equation with the corresponding best network prediction and the absolute error difference. The resulting relative $L^{2}$ error is $2.50\times 10^{-3}$. The hyper-parameter configuration is specified in Section 3.
Table 5: Belousov-Zhabotinsky Equation: Relative $L^{2}$ errors of $u$, $v$ and $w$ equations for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
Ablation Settings Performance: relative $L^{2}$ error
PBC IC1 IC2 SA mini-batch u-equation v-equation w-equation
$3.49\times 10^{-3}$ $2.45\times 10^{-3}$ $2.50\times 10^{-3}$
$6.06\times 10^{-3}$ $4.47\times 10^{-3}$ $4.30\times 10^{-3}$
$7.08\times 10^{-3}$ $4.77\times 10^{-3}$ $4.24\times 10^{-3}$
$5.03\times 10^{-3}$ $3.10\times 10^{-3}$ $3.71\times 10^{-3}$
$9.17\times 10^{-3}$ $9.29\times 10^{-3}$ $8.89\times 10^{-3}$
$1.05\times 10^{-1}$ $1.78\times 10^{-2}$ $1.29\times 10^{-2}$
$2.22\times 10^{-2}$ $1.33\times 10^{-2}$ $1.52\times 10^{-2}$
$1.54\times 10^{-2}$ $9.17\times 10^{-3}$ $1.07\times 10^{-2}$
$9.99\times 10^{-1}$ $9.99\times 10^{-1}$ $9.99\times 10^{-1}$
$7.99\times 10^{-3}$ $5.05\times 10^{-3}$ $5.16\times 10^{-3}$
$9.39\times 10^{-3}$ $9.41\times 10^{-3}$ $9.85\times 10^{-3}$
$3.98\times 10^{-2}$ $2.15\times 10^{-2}$ $2.79\times 10^{-2}$
$4.48\times 10^{-2}$ $2.71\times 10^{-2}$ $3.22\times 10^{-2}$
$7.87\times 10^{-2}$ $3.95\times 10^{-2}$ $5.48\times 10^{-2}$
3.7 Nonlinear Schröedinger Equation
In theoretical physics, nonlinear Schröedinger equation is a nonlinear PDE, applicable to classical and quantum mechanics kato1987nonlinear ; kevrekidis2001discrete ; fibich2015nonlinear . The dimensionless equation of the classical field is

### v3

@@ B.2 Allen-Cahn PDE
The Allen-Cahn type PDE is a classical phase-field model that has been widely used to study phase separation phenomena Bazant (2017). It has numerous practical applications across a range of fields, including materials science Allen & Cahn (1979); Shen & Yang (2010), biological systems Hyman et al. (2014); Takatori & Brady (2015), and electrochemical systems Horstmann et al. (2013); Tian et al. (2015). We begin with the $1$-dim Allen-Cahn equation with periodic boundary conditions formulated as follows:
$u_{t}-\gamma_{1}u_{xx}+\gamma_{2}(u^{3}-u)=0,\,(t,x)\in(0,T]\times(a,b),$ (5)
$u(0,x)=u_{0}(x),\,x\in[a,b],\quad u(t,a)=u(t,b),\,t\in[0,T],$
@@
where $\gamma_{1},\gamma_{2}>0,T>0,a<b$ are prescribed constants. As $\gamma_{2}$ increases, the transition interface of the solutions is sharper, which makes it harder to solve the AC equation numerically.
B.2.1 Case II
We first tested on Allen-Cahn equation with the following, $u_{0}(x)=x^{2}\cos(\pi x)$, $T=1$, a=1a=-1, $b=1$, $\gamma_{1}=0.001$, and $\gamma_{2}=5$.
@@ B.2.1 Case II
For this example, we first test the baseline PINN approach Raissi et al. (2019). As shown in Table 2, the baseline PINN alone fails to accurately solve the Allen-Cahn equation, motivating the development of a more effective training pipeline. To address this, we employ all proposed techniques (except mini-batching) and conduct an ablation study by disabling individual components one at a time. Additionally, we introduce two different decay functions for the initial condition, $e^{-1t}$ and $e^{-0.1t}$, to investigate their influence on the solution. The results are summarized in Table 2. The full scheme, including self-adaptive loss balancing, achieves the best performance, reaching a relative $L^{2}$ error of $8.29\times 10^{-4}$ and relative $L^{1}$ error of $5.34\times 10^{-4}$. In particular, using the faster-decaying initial condition ($IC1$) leads to better results than the slower-decaying case ($IC2$), suggesting that the initial condition plays a more critical role during the early stages of training but becomes less dominant over time. The ablation results further highlight the importance of each methodological component. Removing any single part generally leads to performance degradation, with the most severe impact observed when the initial condition is omitted, resulting in a relative error close to $0.99$ in both $L^{1}$ and $L^{2}$ norms. Figure 4 presents cross-sectional comparisons at different times, further confirming the excellent agreement between the predicted and true solutions across the entire time horizon. These results provide strong evidence for the effectiveness of the proposed training strategy and underscore the critical role of each pipeline component.
Ablation Settings Case II Case IIII
@@
$8.29\times 10^{-4}$ $5.34\times 10^{-4}$ $3.53\times 10^{-4}$ $2.14\times 10^{-4}$
@@
$99\times 10^{-1} 9.9$$9\times 10^{-1} 9.9$$9\times 10^{-1} 9.9$$99\times 10^{-1}$
$99\times 10^{-1} 9.9$$9\times 10^{-1} 9.9$$9\times 10^{-1} 9.9$$99\times 10^{-1}$
@@
$99\times 10^{-1} 9.9$$98\times 10^{-1}$ $5.11\times 10^{-1}$ $3.28\times 10^{-1}$
Table 2: Allen-Cahn (Casse II and Case IIII): Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
B.2.2 Case IIII
For the second case, we change the IC to have highly oscillatory data by $u(0,x)=x^{2}\sin(2\pi x)$. Other parameters are $T=1$, a=1a=-1, $b=1$, $\gamma_{1}=0.001$ and $\gamma_{2}=4$.
@@ B.2.2 Case IIII
Figure 4: Best solution: Allen-Cahn.
B.3 Kuramoto-Sivashinsky PDE
Kuramoto–Sivashinsky (KS) equation is a fourth-order nonlinear partial differential equation, known for its chaotic behavior Kudryashov (1990); Michelson (1986). It has served as a pivotal model in the study of complex dynamical phenomena observed in various physical systems. Consequently, the KS equation provides an excellent case study to demonstrate how PINNs can effectively simulate chaotic dynamics. The equation is given by:
$u_{t}=-u_{xx}-u_{xxxx}-uu_{x},\quad(t,x)\in(0,T]\times(a,b),$ (6)
$u(0,x)=\cos(\frac{x}{16})(1+\sin\frac{x-1}{16}),\quad x\in[a,b],\quad u(t,a)=u(t,b),\,t\in[0,T],$
where $T=20,a=0,b=32\pi$. The best results are shown in 5:
Figure 5: Best solution: Kuramoto-Sivashinsky.
@@ B.3 Kuramoto-Sivashinsky PDE
PBC IC1 IC2 SA Rel. $L^{2}$ error Rel. $L^{1}$ error
@@
Table 3: Kuramoto-Sivashinsky: Relative $L^{2}$ and $L^{1}$ errors for an ablation study illustrating the impact of disabling individual components of the proposed technique and training pipeline.
As in Table 3, solving this PDE system using a baseline PINN alone proves to be exceptionally challenging. The incorporation of periodic neural networks for boundary conditions, paired with initial condition IC2, consistently yields superior results compared to IC1. This suggests that, in contrast to the Allen-Cahn equation discussed in Section B.2, the choice of initial condition plays a more significant role in the training process for this more complex system. This observation underscores the importance of carefully selecting both the boundary conditions and initial conditions when training neural networks for chaotic or higher-order PDE systems, where these factors contribute notably to the accuracy and stability of the solution.
B.4 Cahn-Hilliard Equation
The Cahn-Hilliard equation, a variant of the Allen-Cahn equation, also models phase separation. This process describes how the components of a binary fluid spontaneously separate into distinct domains, each composed of a single component Cahn & Hilliard (1958); Kim et al. (2016); Lee et al. (2014); Miranville (2017). The equation is expressed as:
$u(0,x)=u_{0}(x),\quad x\in[-L,L],$
$u(t,-L)=u(t,L),\quad t\in[0,T],$
Figure 6: Best solution of the Cahn-Hilliard. The resulting relative $L^{2}$ error is $3.53\times 10^{-4}$.
In this experiment, the full scheme yields the best results. Additionally, methods using initial condition IC2 outperform those with IC1. Except for the full schemes and the experiment with periodic neural networks for boundary conditions and initial condition IC2, none of the other approaches produce accurate solutions. The errors for all experiments are summarized in Table 4. The best results are illustrated in Figure 6.
Ablation Settings Performance
PBC IC1 IC2 SA Rel. $L^{2}$ error Rel. $L^{1}$ error
$1.06\times 10^{-3}$ $5.61\times 10^{-4}$
$9.75\times 10^{-4}$ $5.65\times 10^{-4}$
$2.50\times 10^{-1}$ $5.87\times 10^{-2}$
$3.87\times 10^{-3}$ $1.91\times 10^{-3}$
$99\times 10^{-1} 9.9$$99\times 10^{-1}$
$99\times 10^{-1} 9.9$$99\times 10^{-1}$
$3.48\times 10^{-1}$ $2.94\times 10^{-1}$
$2.37\times 10^{-1}$ $1.09\times 10^{-1}$
$4.07\times 10^{-1}$ $3.45\times 10^{-1}$
$1.55\times 10^{-2}$ $8.20\times 10^{-3}$
$99\times 10^{-1} 9.9$$99\times 10^{-1}$
$99\times 10^{-1} 9.9$$98\times 10^{-1}$
B.5 Gray-Scott Equation
Reaction and diffusion of chemical species can produce a variety of patterns (De Kepper et al., 1991; Nishiura & Ueyama, 1999), reminiscent of those often seen in nature. The Gray-Scott type system is one of classical mathematical models for chemical reactions (Gray & Scott, 1983, 1984; Liang et al., 2022). The general irreversible Gray-Scott equations describe such reactions:
@@ B.5 Gray-Scott Equation
$u_{t}=\epsilon_{1}u_{xx}+b(1-u)-uv^{2},$ (9)
$v_{t}=\epsilon_{2}v_{xx}-(b+k)v+uv^{2},\quad(t,x)\in(0,T]\times(-L,L),$
$u(0,x)=u_{0}(x),\quad v(0,x)=v_{0}(x),\quad x\in[-L,L],$
$u(t,-L)=u(t,L),\quad v(t,-L)=v(t,L),\quad t\in[0,T],$
where $T=20,L=50$, $\epsilon_{1}=1,\epsilon_{2}=0.01$ are diffusion rates, $b=0.02$ is the ”feeding rate” that adds $U$, $k=0.0562$ is the ”killing rate” that removes $V$. We set our initial conditions as:
$u_{0}(x)=1-\frac{\sin(\pi(x-50)/100)^{4}}{2},$
$v_{0}(x)=\frac{\sin(\pi(x-50)/100)^{4}}{4}.$
In this ablation study, we not only evaluate the performance of self-adaptive neural networks but also test our scheme with mini-batching on the Gray-Scott equations, which model the densities of two interacting species. For mini-batching, we use 5 batches, each containing 3,280 samples. As shown in Table 5, the results confirm that each proposed component improves the overall model performance. Omitting any one of these components leads to higher error rates. Notably, the errors for the $u$-equation are typically smaller than those for the $v$-equation, which aligns with our expectations. This is because the $v$-equation is sharper and more challenging to train compared to the $u$-equation. Additionally, experiments using initial condition IC2 consistently outperform those using IC1. Specifically, the full scheme with IC2 achieves the best results, with a relative $L^{2}$ error of $2.05\times 10^{-4}$ for $u$-equation and $8.46\times 10^{-4}$ for $v$-equation. These results suggest that IC2 is more effective for larger time-scale PDEs, as IC1 causes the initial condition to decay too quickly, reducing its influence during training. The predicted solutions from our best model are visualized in Figures 7 and 8, respectively.
Figure 7: Best solution of the Gray-Scott (u species). The resulting relative $L^{2}$ error is $2.05\times 10^{-4}$.
Figure 8: Best solution of the Gray-Scott (v species). The resulting relative $L^{2}$ error is $8.46\times 10^{-4}$.
@@
PBC IC1 IC2 SA mini-batch $\mathcal{E}_{2}:u$ $\mathcal{E}_{2}:v$
@@
B.6 Nonlinear Schröedinger Equation
In theoretical physics, nonlinear Schröedinger equation is a nonlinear PDE, applicable to classical and quantum mechanics (Kato, 1987; Kevrekidis et al., 2001; Fibich, 2015). The dimensionless equation of the classical field is